OpenStudy (anj123):

Find the area of the region enclosed by the inner loop of the curve. r = 2 + 4 sin(θ) The answer I got was -4sqrt(3) but thats wrong.

satellite73 (satellite73):

unlikely that the area is negative...

satellite73 (satellite73):

i totally forget how to do this, but i think your main step is to find the limits of integration

satellite73 (satellite73):

for the inner loop, start where \(r=0\) so \[\theta=\frac{7\pi}{6}\](i think)